Please solve this using the elimination method.
2x + 3y = 2
-28 = 4x - 2y

Answer by checkley77(12844)

2x+3y=2 multiply by 2 & add.
-4x+2y=28
4x+6y=4
----------------
8y=32
y=32/8
y=4 ans.
2x+3*4=2
2x+12=2
2x=2-12
2x=-10
x=-10/2
x=-5 ans.
Proof:
-4*-5+2*4=28
20+8=28
28=28